OpenStudy (anonymous):

Solve this problem in your notebook using all four steps. Harvey is 3 times as old as Jane. The sum of their ages is 48 years. Find the age of each. Jane is a0 years old. Harvey is a1 years old.

OpenStudy (anonymous):

let h = H's age let j = J's age : Write an equation for each statement: ; "Harvey is 3 times as old as Jane." h = 3j : "The sum of their ages is 48 years." h + j = 48 From the 1st statement, replace h with 3j 3j + j = 48 4j = 48 j = 48/4 j = 12 yrs is J's age then 3(12) = 36 yrs is H's age